OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x – 5y + 2z = 7 4x + 9y – z = 2 x – 3y + 3z = 7 11x + 13y = 11 13x + 24y = 13 11x + 13y = 11 3x – 5y = 7 x – 3y = 7 13x + 24y = 13 x – 3y = 7 3x – 5y = 7

OpenStudy (mysesshou):

(eq1) 3x – 5y + 2z = 7 (eq2) 4x + 9y – z = 2 (eq3) x – 3y + 3z = 7 (eq1) 3x – 5y + 2z = 7 (eq2) 4x + 9y – z = 2 multiply whole eq2 by 2 gives eq2.1 (eq2.1) 8x + 18y -2z = 4 add eq2.1 to eq1 So: (eq1) 3x – 5y + 2z = 7 (eq2.1) 8x + 18y -2z = 4 = 11x + 13y = 11 <---- (eq2) 4x + 9y – z = 2 (eq3) x – 3y + 3z = 7 next, multiply all of eq2 by 3, giving eq2.2 (eq2.2) 12x + 27y – 3z = 6 Then add eq2.2 and eq3 together (eq2.2) 12x + 27y – 3z = 6 (eq3) x – 3y + 3z = 7 =13x + 24y = 13 <----- which is the first pair.

OpenStudy (mysesshou):

Hope someone proofs this ... i think it is right though. :)

OpenStudy (mysesshou):

this is elimination using the eq2 for each one, and then eq2 is eliminated

OpenStudy (anonymous):

thank you

OpenStudy (mysesshou):

when you do it in the future though, make sure you only multiply through using one equation, and then that is the equation removed